# Meta-automatic Sequences

### v1

@@ 2.1 Denestability
and then set a(2n+1)=1a(2n)a(2n+1)=1-a(2n) to preserve (4). Under balance, however, both schemes collapse to ultimate periodicity, and this is formalized below.
@@
    If $a(2n)=a(n+a(n))$ for all $n\geq 0$, then $a(0)=0$ and
    $a(2n)=0,\qquad a(2n+1)=1\qquad(n\geq 2),$
    so $a$ is ultimately periodic of period $2$.
@@
    If a(2n)=a(n+1a(n))a(2n)=a(n+1-a(n)) for all $n\geq 0$, then $a(0)=1$ and
    $a(2n)=1,\qquad a(2n+1)=0\qquad(n\geq 2),$
    so $a$ is ultimately periodic of period $2$.
@@
(a) Taking $n=0$ gives $a(0)=a(a(0))$. If $a(0)=1$, then balancedness forces $a(1)=0$, and hence $a(0)=a(1)=0$, a contradiction. Thus $a(0)=0$ and $a(1)=1$.
Write $b_{m}=a(2m)$, so that balancedness gives $a(2m+1)=1-b_{m}$. Using $a(2n)=a(n+a(n))$ with $n=2m$, we obtain
$b_{2m}=a(4m)=a(2m+a(2m))=a(2m+b_{m})=\begin{cases}a(2m)=b_{m},&b_{m}=0;\\[2.0pt] a(2m+1)=1-b_{m}=0,&b_{m}=1.\end{cases}$
@@
Next, using $a(2n)=a(n+a(n))$ with $n=2m+1$ and $a(2m+1)=1-b_{m}$, we get
$b_{2m+1}=a(4m+2)=a(2m+1+a(2m+1))=a(2m+2-b_{m})=\\ \begin{cases}a(2m+2)=b_{m+1},&b_{m}=0;\\[2.0pt] a(2m+1)=0,&b_{m}=1.\end{cases}$
@@
Lemma 14 shows that meta-automatic behavior requires passing to base 44: the dyadic pairing (4) is preserved, but there is now room for selectors one scale below.
2.2 A non-nested baseline case
Fix $a(0)=0$, $a(1)=1$, and define
@@ 2.2 A non-nested baseline case
so that $a(2n)+a(2n+1)=1$ holds automatically. The three main sequences investigated in this paper arise by specializing (5). To begin with, we consider an integer sequence $(\mathcal{Q}(n))_{n\geq 0}$ that does not have a nested structure in terms of its recurrence but that is to be used extensively in our subsequent work.
@@
$\mathcal{Q}(4n)$ $=\mathcal{Q}(n),$ $\mathcal{Q}(4n+1)$ $=1-\mathcal{Q}(n),$ (6)
$\mathcal{Q}(4n+2)$ $=\mathcal{Q}(2n),$ $\mathcal{Q}(4n+3)$ $=1-\mathcal{Q}(2n).$ (7)
Omitting details for brevity, we have that there is a $3$-state DFAO for $(\mathcal{Q}(n))_{n\geq 0}$, in such a way so that $\mathcal{Q}$ is $4$-automatic and computed by the DFAO with states $\{A,B,C\}$, initial state $A$, and output $\tau(A)=\tau(B)=0$, $\tau(C)=1$. This DFAO is shown in Figure 1. The DFAO for $\mathcal{Q}$ corresponds to the $4$-uniform morphism
@@
with coding $\tau(A)=\tau(B)=0$, $\tau(C)=1$.
@@
Define $d:\mathbb{N}_{0}\to\{0,1\}$ by $d(0)=0$ and $d(n)=\lfloor\log_{2}n\rfloor\bmod 2$ for $n\geq 1$. Then
@@
3 Meta(1)
Definition 17 (Meta(1)).
Define $\mathcal{M}_{1}:\mathbb{N}_{0}\to\{0,1\}$ by $\mathcal{M}_{1}(0)=0$, $\mathcal{M}_{1}(1)=1$, and for $n\geq 0$, so that
$\mathcal{M}_{1}(4n)=\mathcal{M}_{1}(n),\ \mathcal{M}_{1}(4n+1)=1-\mathcal{M}_{1}(n),$ (8)
and
$\mathcal{M}_{1}(4n+2)=\mathcal{M}_{1}(2n+1-\mathcal{M}_{1}(n)),\ \mathcal{M}_{1}(4n+3)=1-\mathcal{M}_{1}(2n+1-\mathcal{M}_{1}(n)).$ (9)
@@ 3 Meta(1)
Well-definedness follows by induction on $m$. For $m=4n+r$, the right-hand sides in (8) reduce to indices <m<m; for (9), the selector index is $2n+1-\mathcal{M}_{1}(n)\in\{2n,2n+1\}<4n+2\leq m$. Balancedness follows from (5), with $\mathcal{M}_{1}(4n+1)=1-\mathcal{M}_{1}(4n)$ and $\mathcal{M}_{1}(4n+3)=1-\mathcal{M}_{1}(4n+2)$. ∎
@@
Because $\mathcal{M}_{1}$ is balanced, Lemma 10 yields
@@
so (9) is equivalent to the non-nested relation $\mathcal{M}_{1}(4n+2)=\mathcal{M}_{1}(2n+1)\oplus\mathcal{M}_{1}(n)$. The selector form is the natural “meta” description and is the form that generalizes.
@@ 3.1 A 4-state DFAO
Define the auxiliary function
$b(n):=\mathcal{M}_{1}(2n+1-\mathcal{M}_{1}(n)).$ (10)
Consider the lifted state $s(n):=(\mathcal{M}_{1}(n),b(n))\in\{0,1\}^{2}$.
@@
For each $r\in\{0,1,2,3\}$, the state $s(4n+r)$ depends only on $s(n)$.
@@
Write $s(n)=(x,y)$ where $x=\mathcal{M}_{1}(n)$ and $y=b(n)$. By the balancedness given by Lemma 18 and the XOR linearization (Lemma 10), we have $y=b(n)=\mathcal{M}_{1}(2n+1)\oplus\mathcal{M}_{1}(n)=(1-\mathcal{M}_{1}(2n))\oplus x$, so that
$\mathcal{M}_{1}(2n)=1\oplus x\oplus y,\qquad\mathcal{M}_{1}(2n+1)=x\oplus y.$ (11)
Case $r=0$. The first coordinate is $\mathcal{M}_{1}(4n)=\mathcal{M}_{1}(n)=x$. For the second coordinate, $b(4n)=\mathcal{M}_{1}(8n+1-\mathcal{M}_{1}(4n))=\mathcal{M}_{1}(8n+1-x)$. If $x=0$, then $b(4n)=\mathcal{M}_{1}(8n+1)=\mathcal{M}_{1}(4(2n)+1)=1-\mathcal{M}_{1}(2n)=1-(1\oplus y)=y$. If $x=1$, then $b(4n)=\mathcal{M}_{1}(8n)=\mathcal{M}_{1}(4(2n))=\mathcal{M}_{1}(2n)=1\oplus x\oplus y=y$. In both cases, we find that s(4n)=(x,y)s(4n)=(x,\,y).
Case $r=1$. The first coordinate is $\mathcal{M}_{1}(4n+1)=1-x=1\oplus x$. For the second coordinate, we have that $b(4n+1)=\mathcal{M}_{1}(8n+2+x)$. If $x=0$, then $b(4n+1)=\mathcal{M}_{1}(8n+2)=\mathcal{M}_{1}(4(2n)+2)=b(2n)$. Using the property that $b(2n)=\mathcal{M}_{1}(4n+1-\mathcal{M}_{1}(2n))$, together with $\mathcal{M}_{1}(2n)=1\oplus y$, this yields $\mathcal{M}_{1}(4n+y)$. So, if $y=0$, then $\mathcal{M}_{1}(4n)=x=0$, and if $y=1$, then $\mathcal{M}_{1}(4n+1)=1-x=1$, so that $b(4n+1)=y$. Similarly, if $x=1$, then $b(4n+1)=\mathcal{M}_{1}(8n+3)=\mathcal{M}_{1}(4(2n)+3)=1-b(2n)$. With $\mathcal{M}_{1}(2n)=y$, $b(2n)=\mathcal{M}_{1}(4n+1-y)$, we find that if $y=0$, then $1-\mathcal{M}_{1}(4n+1)=x=1$, and if $y=1$, then $1-\mathcal{M}_{1}(4n)=1-x=0$, so that $b(4n+1)=1\oplus y$. Consequently, we obtain that $s(4n+1)=(1\oplus x,\;x\oplus y)$.
Case $r=2$. The first coordinate is $\mathcal{M}_{1}(4n+2)=b(n)=y$. For the second coordinate, $b(4n+2)=\mathcal{M}_{1}(8n+5-y)$. If $y=0$, then $\mathcal{M}_{1}(8n+5)=\mathcal{M}_{1}(4(2n+1)+1)=1-\mathcal{M}_{1}(2n+1)=1\oplus(x\oplus y)=1\oplus x$. If $y=1$, then $\mathcal{M}_{1}(8n+4)=\mathcal{M}_{1}(4(2n+1))=\mathcal{M}_{1}(2n+1)=x\oplus y=1\oplus x$. In both cases, we see that $s(4n+2)=(y,\;1\oplus x)$.
Case $r=3$. The first coordinate is $\mathcal{M}_{1}(4n+3)=1-y=1\oplus y$. For the second coordinate, $b(4n+3)=\mathcal{M}_{1}(8n+6+y)$. If $y=0$, then $\mathcal{M}_{1}(8n+6)=\mathcal{M}_{1}(4(2n+1)+2)=b(2n+1)$. Since $\mathcal{M}_{1}(2n+1)=x$, $b(2n+1)=\mathcal{M}_{1}(4n+3-x)$. So, if $x=0$, then $\mathcal{M}_{1}(4n+3)=1\oplus y=1$; if $x=1$, then $\mathcal{M}_{1}(4n+2)=y=0$; so $b(4n+3)=x\oplus y\oplus 1$. If $y=1$, then $\mathcal{M}_{1}(8n+7)=\mathcal{M}_{1}(4(2n+1)+3)=1-b(2n+1)$. Since $\mathcal{M}_{1}(2n+1)=x\oplus 1$, we find that $b(2n+1)=\mathcal{M}_{1}(4n+2+x)$. So, if $x=0$, then $1-\mathcal{M}_{1}(4n+2)=1-y=0$; if $x=1$, then $1-\mathcal{M}_{1}(4n+3)=y=1$, so that $b(4n+3)=x\oplus y\oplus 1$. Thus $s(4n+3)=(1\oplus y,\;x\oplus y\oplus 1)$. ∎
Theorem 21 (DFAO for $\mathcal{M}_{1}$).
@@
The DFAO of $\mathcal{M}_{1}$ has $\mathbb{F}_{2}$-affine transitions, as does that of $\mathcal{M}_{2}$ (Section 4). All three sequences $\mathcal{Q}$, $\mathcal{M}_{1}$, and $\mathcal{M}_{2}$ arise from the same mechanism: The balance constraint turns meta-automatic selectors into $\mathbb{F}_{2}$-affine maps on the lifted state space.
Theorem 23.
Neither $\mathcal{M}_{1}$ nor $\mathcal{M}_{2}$ is denestable: In each case, the closure of the $4$-kernel over $\mathbb{Z}$ requires a lifted state of dimension $2$. Moreover, both sequences are $\mathbb{F}_{2}$-linearizable, with
$\mathcal{M}_{1}(4n+2)=\mathcal{M}_{1}(2n+1)\oplus\mathcal{M}_{1}(n),$
for $\mathcal{M}_{1}$, and
Proof.
The $\mathbb{F}_{2}$-linearization follows directly from Lemma 10: Since $\mathcal{M}_{1}$ is balanced, we see that $\mathcal{M}_{1}(2n+1-\mathcal{M}_{1}(n))=\mathcal{M}_{1}(2n+1)\oplus\mathcal{M}_{1}(n)$, giving a fixed-argument recurrence.
For non-denestability, suppose the $4$-kernel could be closed using $\mathcal{M}_{1}(n)$ alone, without a second state component. The auxiliary sequence $b(n)=\mathcal{M}_{1}(2n+1-\mathcal{M}_{1}(n))$ equals $\mathcal{M}_{1}(2n)$ when $\mathcal{M}_{1}(n)=1$ and $\mathcal{M}_{1}(2n+1)$ when $\mathcal{M}_{1}(n)=0$. Since $\mathcal{M}_{1}(2n)\neq\mathcal{M}_{1}(2n+1)$ by balance, no single fixed argument $cn+d$ over $\mathbb{Z}$ reproduces both cases, as $b$ is not of the form $\mathcal{M}_{1}(cn+d)$. Thus the $4$-kernel cannot be closed over $\mathbb{Z}$ without the lifted state, confirming that $\mathcal{M}_{1}$ is not denestable. ∎
@@
This stands in contrast to Allouche and Shallit [4], who show that the frequency sequence $F(n)$ of the Hofstadter variant $V=Q_{1,4}$ satisfies a recurrence of the form F(2a)=g(F(a2),F(a1),F(a),F(a+1))F(2a)=g(F(a-2),F(a-1),F(a),F(a+1)) with fixed shifts, making $F$ directly denestable in our sense (Definition 12). In our setting, the sequences themselves (not derived quantities) are meta-automatic from the start, and it is the balance constraint that recovers automaticity via $\mathbb{F}_{2}$-linearization rather than denestability.
Corollary 25 (Uniform morphism for $\mathcal{M}_{1}$).

### v2

@@ 2.1 Base-4 recurrences
and we let a(2n+1)=1a(2n)a(2n+1)=1-a(2n). In both cases, we obtain ultimate periodicity, as we now show.
@@
    If $a(2n)=a(n+a(n))$ for all $n\geq 0$, then $a(0)=0$ and $a$ is ultimately $2$-periodic, with
    $a(2n)=0,\qquad a(2n+1)=1\qquad(n\geq 2).$
@@
    If a(2n)=a(n+1a(n))a(2n)=a(n+1-a(n)) for all $n\geq 0$, then $a(0)=1$ and $a$ is ultimately $2$-periodic, with
    $a(2n)=1,\qquad a(2n+1)=0\qquad(n\geq 2).$
@@
(a) Taking $n=0$ gives $a(0)=a(a(0))$. If $a(0)=1$, then balancedness gives us that $a(1)=0$, and hence $a(0)=a(1)=0$, a contradiction. Thus $a(0)=0$ and $a(1)=1$.
Write $b_{m}=a(2m)$, so that balancedness gives that $a(2m+1)=1-b_{m}$. Using $a(2n)=a(n+a(n))$ with $n=2m$, we obtain
$b_{2m}=a(4m)=a(2m+a(2m))=a(2m+b_{m})=\begin{cases}a(2m)=b_{m},&\text{if }b_{m}=0;\\ a(2m+1)=1-b_{m}=0,&\text{if }b_{m}=1.\end{cases}$
@@
Now, using $a(2n)=a(n+a(n))$ with $n=2m+1$ and $a(2m+1)=1-b_{m}$, we get
$b_{2m+1}=a(4m+2)=a(2m+1+a(2m+1))=a(2m+2-b_{m})=\\ \begin{cases}a(2m+2)=b_{m+1},&\text{if }b_{m}=0;\\ a(2m+1)=0,&\text{if }b_{m}=1.\end{cases}$
@@
Lemma 6 leads us to consider base-44 recurrences such that the balance identity in (3) is preserved.
2.2 A non-nested case
Fix $a(0)=0$ and $a(1)=1$, and define
@@ 2.2 A non-nested case
so that $a(2n)+a(2n+1)=1$ holds automatically. The sequences studied in this paper arise as specializations of (5). We first consider a sequence $(\mathcal{Q}(n))_{n\geq 0}$ whose recurrence is not nested, but which is used in the sequel.
@@
$\mathcal{Q}(4n)$ $=\mathcal{Q}(n),$ $\mathcal{Q}(4n+1)$ $=1-\mathcal{Q}(n),$
$\mathcal{Q}(4n+2)$ $=\mathcal{Q}(2n),$ $\mathcal{Q}(4n+3)$ $=1-\mathcal{Q}(2n).$
The sequence $\mathcal{Q}$ is $4$-automatic. It is computed by the $3$-state DFAO with states in $\{A,B,C\}$, initial state $A$, and output $\tau(A)=\tau(B)=0$, $\tau(C)=1$, shown in Figure 1. This DFAO corresponds to the $4$-uniform morphism
@@
along with a coding such that $\tau(A)=\tau(B)=0$ and $\tau(C)=1$.
@@
Define $d:\mathbb{N}_{0}\to\{0,1\}$ so that $d(0)=0$ and $d(n)=\lfloor\log_{2}n\rfloor\bmod 2$ for $n\geq 1$. Then
@@
3 The sequence $\mathcal{M}_{1}$
Definition 9.
Define $\mathcal{M}_{1}(0)=0$ and $\mathcal{M}_{1}(1)=1$. For $n\geq 1$, set
$\mathcal{M}_{1}(4n)=\mathcal{M}_{1}(n),\qquad\mathcal{M}_{1}(4n+1)=1-\mathcal{M}_{1}(n).$ (6)
For $n\geq 0$, set
$\mathcal{M}_{1}(4n+2)=\mathcal{M}_{1}(2n+1-\mathcal{M}_{1}(n)),\qquad\mathcal{M}_{1}(4n+3)=1-\mathcal{M}_{1}(2n+1-\mathcal{M}_{1}(n)).$ (7)
@@ 3 The sequence $\mathcal{M}_{1}$
Well-definedness follows by induction on $m$. For $m=4n+r$, the right-hand sides in (6) reduce to indices <m<m; for (7), the nested argument $2n+1-\mathcal{M}_{1}(n)\in\{2n,2n+1\}<4n+2\leq m$. Balancedness follows from (5), since $\mathcal{M}_{1}(4n+1)=1-\mathcal{M}_{1}(4n)$ and $\mathcal{M}_{1}(4n+3)=1-\mathcal{M}_{1}(4n+2)$. ∎
@@
Since $\mathcal{M}_{1}$ is balanced, Lemma 5 gives
@@
so (7) is equivalent to the non-nested relation $\mathcal{M}_{1}(4n+2)=\mathcal{M}_{1}(2n+1)\oplus\mathcal{M}_{1}(n)$.
@@ 3.1 A 4-state DFAO
Define
$b(n):=\mathcal{M}_{1}(2n+1-\mathcal{M}_{1}(n)),$ (8)
and consider the pair $s(n):=(\mathcal{M}_{1}(n),b(n))$. Informally, the following result gives us that the pair $s(4n+r)$ depends only on $s(n)$, for $r\in\{0,1,2,3\}$ and for $n\geq 1$, disregarding the trivial $n=0$ case whereby $\mathcal{M}_{1}(0)=0$.
@@
Let $x=\mathcal{M}_{1}(n)$ and $y=b(n)$, and let $n\geq 1$. Then
$s(4n)$ $=(x,y),$ (9)
$s(4n+1)$ $=(1\oplus x,\;x\oplus y),$ (10)
$s(4n+2)$ $=(y,\;1\oplus x),\text{and}$ (11)
$s(4n+3)$ $=(1\oplus y,\;x\oplus y\oplus 1).$ (12)
@@
By Lemma 10 and Lemma 5, we have $y=b(n)=\mathcal{M}_{1}(2n+1)\oplus\mathcal{M}_{1}(n)=(1-\mathcal{M}_{1}(2n))\oplus x$, so that
$\mathcal{M}_{1}(2n)=1\oplus x\oplus y,\qquad\mathcal{M}_{1}(2n+1)=x\oplus y.$
Case $r=0$. The first coordinate is $\mathcal{M}_{1}(4n)=\mathcal{M}_{1}(n)=x$. The latter coordinate satisfies $b(4n)=\mathcal{M}_{1}(8n+1-\mathcal{M}_{1}(4n))=\mathcal{M}_{1}(8n+1-x)$. If $x=0$, then $b(4n)=\mathcal{M}_{1}(8n+1)=\mathcal{M}_{1}(4(2n)+1)=1-\mathcal{M}_{1}(2n)=1-(1\oplus x\oplus y)=1-(1\oplus y)=y$. If $x=1$, then $b(4n)=\mathcal{M}_{1}(8n)=\mathcal{M}_{1}(4(2n))=\mathcal{M}_{1}(2n)=1\oplus x\oplus y=y$. In both cases, we find that (9) holds.
Case $r=1$. The first coordinate is $\mathcal{M}_{1}(4n+1)=1-x=1\oplus x$. The latter coordinate satisfies $b(4n+1)=\mathcal{M}_{1}(8n+2+x)$. If $x=0$, then $b(4n+1)=\mathcal{M}_{1}(8n+2)=\mathcal{M}_{1}(4(2n)+2)=b(2n)$. Using the property that $b(2n)=\mathcal{M}_{1}(4n+1-\mathcal{M}_{1}(2n))$, together with $\mathcal{M}_{1}(2n)=1\oplus y$, this yields $b(2n)=\mathcal{M}_{1}(4n+y)$. So, if $y=0$, then $\mathcal{M}_{1}(4n)=x=0$, and if $y=1$, then $\mathcal{M}_{1}(4n+1)=1-x=1$, so that $b(4n+1)=y$. Similarly, if $x=1$, then $b(4n+1)=\mathcal{M}_{1}(8n+3)=\mathcal{M}_{1}(4(2n)+3)=1-b(2n)$. With $\mathcal{M}_{1}(2n)=y$, $b(2n)=\mathcal{M}_{1}(4n+1-y)$, we find that if $y=0$, then $1-\mathcal{M}_{1}(4n+1)=x=1$, and if $y=1$, then $1-\mathcal{M}_{1}(4n)=1-x=0$, so that $b(4n+1)=1\oplus y$. Consequently, we obtain that (10) holds.
Case $r=2$. The first coordinate is $\mathcal{M}_{1}(4n+2)=b(n)=y$. The latter coordinate satisfies $b(4n+2)=\mathcal{M}_{1}(8n+5-y)$. If $y=0$, then $\mathcal{M}_{1}(8n+5)=\mathcal{M}_{1}(4(2n+1)+1)=1-\mathcal{M}_{1}(2n+1)=1\oplus(x\oplus y)=1\oplus x$. If $y=1$, then $\mathcal{M}_{1}(8n+4)=\mathcal{M}_{1}(4(2n+1))=\mathcal{M}_{1}(2n+1)=x\oplus y=1\oplus x$. In both cases, we see that (11) holds.
Case $r=3$. The first coordinate is $\mathcal{M}_{1}(4n+3)=1-y=1\oplus y$. The latter coordinate satisfies $b(4n+3)=\mathcal{M}_{1}(8n+6+y)$. If $y=0$, then $\mathcal{M}_{1}(8n+6)=\mathcal{M}_{1}(4(2n+1)+2)=b(2n+1)$. Since $\mathcal{M}_{1}(2n+1)=x$, $b(2n+1)=\mathcal{M}_{1}(4n+3-x)$. So, if $x=0$, then $\mathcal{M}_{1}(4n+3)=1\oplus y=1$; if $x=1$, then $\mathcal{M}_{1}(4n+2)=y=0$; so $b(4n+3)=x\oplus y\oplus 1$. If $y=1$, then $\mathcal{M}_{1}(8n+7)=\mathcal{M}_{1}(4(2n+1)+3)=1-b(2n+1)$. Since $\mathcal{M}_{1}(2n+1)=x\oplus 1$, we find that $b(2n+1)=\mathcal{M}_{1}(4n+2+x)$. So, if $x=0$, then $1-\mathcal{M}_{1}(4n+2)=1-y=0$; if $x=1$, then $1-\mathcal{M}_{1}(4n+3)=y=1$, so that $b(4n+3)=x\oplus y\oplus 1$. This gives us the desired relation in (12). ∎
Theorem 13.
@@
The four transition rules of the DFAO of $\mathcal{M}_{1}$ are $\mathbb{F}_{2}$-affine in $(x,y)$; the same is true for $\mathcal{M}_{2}$ (Section 4). For each of the sequences $\mathcal{Q}$, $\mathcal{M}_{1}$, and $\mathcal{M}_{2}$, the affine form is a consequence of Lemma 5.
The defining recurrence of $\mathcal{M}_{1}$ is not directly an instance of Theorem 2, since the argument $2n+1-\mathcal{M}_{1}(n)$ in the nested recurrence in (7) depends on the value $\mathcal{M}_{1}(n)$. The $b$-sequence defined in (8) removes this dependence. Furthermore, Lemma 5 gives us that
$\mathcal{M}_{1}(4n+2)=\mathcal{M}_{1}(2n+1)\oplus\mathcal{M}_{1}(n).$
The XOR identity is a direct consequence of Lemma 5, since $\mathcal{M}_{1}$ is balanced.
@@
By contrast, Allouche and Shallit [4] show that the frequency sequence $F(n)$ of the Hofstadter variant $V=Q_{1,4}$ satisfies a recurrence of the form F(2a)=g(F(a2),F(a1),F(a),F(a+1))F(2a)=g(F(a-2),F(a-1),F(a),F(a+1)) with fixed shifts. This satisfies the required conditions in Theorem 2. In our setting, the sequences we consider themselves, as opposed to associated frequency sequences, are meta-automatic. In this direction, Lemma 5 provides a key tool, for our purposes.
Corollary 16.